OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. y = 3x + 2 y = x + 12 A. (0, 12) B. (3, 2) C. (5, 17) D. (0, 2)

OpenStudy (shadowlegendx):

Hey!

OpenStudy (shadowlegendx):

Do you know what the substitution method is?

OpenStudy (anonymous):

eh kind of.... But if you could refresh that would be great! :D

OpenStudy (shadowlegendx):

Umm, I'll take that as a no It's basically solving for a variable, then inputting what you got for it. y = 3x + 2 y = x + 12 We already have y solved for, so lets input in for y 3x + 2 = x + 12 Subtract 2 from both sides 3x = x + 10 Subtract x from both sides 2x = 10 Divide both sides by 2 x = 5

OpenStudy (shadowlegendx):

Notice how they both have y = ______ . That means that both equations equal y. Thus meaning that both equations equal each other. So we can put them as equaling each other to solve for x.

OpenStudy (shadowlegendx):

So now we have x = 5 Lets put in 5 for x and solve for y. y = x + 12 y = 5 + 12 y = 17

OpenStudy (shadowlegendx):

Now lets CHECK our ANSWER by inputting in our answers to see if we get it correct. Remember x = 5 y = 17 y = 3x + 2 17 = 3(5) + 2 17 = 15 + 2 17 = 17 That's correct, now lets check the other one. y = x + 12 17 = 5 + 12 17 = 17 That's correct too! So now we can so for sure that: x = 5 and y = 17

OpenStudy (anonymous):

Ok so then it would be (5,17)? So C would be the answer correct?

OpenStudy (shadowlegendx):

Yep. Remember how you solve for slope and all that. Coordinates are put into the format of (x,y) So yeah, (5,17) is correct

OpenStudy (shadowlegendx):

If you need any further help with anything, just link me by posting @ShadowLegendX in your question ;P

OpenStudy (anonymous):

Ok!!! For sure!